OpenStudy (cheddarskittles):

Show work? 5(P+3)+9=3(p-2)+12

OpenStudy (anonymous):

Is this solving for P?

OpenStudy (anonymous):

Eh.. I'm not sure what to solve for then xD

OpenStudy (anonymous):

ok... i think we went through this... 5(p+3)+9=3(p-2)+12 Multiply 5*p=5p and 5*3=15 then 3*p=3p and 3*-2=-6 <----remember that negatives and positives multiplied will change to the negative sign. so youll end up with 5p+15+9=3p-6+12 add the extra numbers that doesnt have a P with it. so youll end up with 5p+24=3p+6 subtract 6 from 24 and end up with 5p+18=3p subtract 5p by 3p and youll get 18=-2p divide -2 on both sides to cancel out the -2 from the p so p will be by itself and the result is -6=p